An experimenter on a ladder releases a marble from rest. the marble takes 4 seconds to fall to the ground. How high was the marble when it was released?
Answer by Alan3354(69443)

An experimenter on a ladder releases a marble from rest. the marble takes 4 seconds to fall to the ground. How high was the marble when it was released?
---------------
If it falls due go gravity on Earth:
32 ft/sec/sec is commonly used for acceleration due to gravity.
--> s = at^2/2 = 16t^2
s = 16*16 = 256 feet
A long ladder.